A Complete Bibliography of Publications in the Journal of Fish Biology (1960–1969)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.01

Title word cross-reference

Account [8]. aculeatus [12, 29]. Acute [1]. Adult [11]. adunconucleata [20]. Age [33]. Albova [20]. American [3]. Ammocoetes [13]. Ammonia [1]. Anadromous [13]. Anguilla [23]. Aquatic [6]. Atlantic [28]. auratus [4]. Autumn [31]. Avon [18].

Bass [3]. Batsch [18]. Beds [31]. Between [14, 30]. Bibliography [17]. Biology [18]. Blood [32]. Book [10, 15, 25, 35]. borealis [20]. Brackish [22]. Breeding [22]. British [8, 3, 27]. Brook [9]. Brown [30, 14].

Carassius [4]. Carbon [32]. Carp [6]. cephalus [9]. Cestode [18]. Changes [28]. Chub [9]. Cilan [24]. Ciliata [20]. Comparison [13]. Concentrations [1]. Control [6].

Dace [9, 18, 21, 2]. Dead [22]. Determination [33]. Development [13, 21].

Diagnosis [28]. Dioxide [32]. Distribution [8]. Dogiel [20]. Dyrnogydd [24].

Early [21]. Ecology [22]. Eels [23]. Effects [32]. Egg [30]. England [7]. Enlarger [5]. Erne [26]. Esox [19].

Fauna [12, 29]. Fecundity [14, 19]. Feeding [7, 23]. Field [34]. Film [16]. Fish [5, 29, 34, 22, 8, 7, 17, 24]. Fishes [27]. Fluctuating [1]. Food [14, 2]. Forms [13]. Freshwater [8]. Fry [30]. Furunculosis [4].

gairdneri [32]. Gasterosteus [12, 29]. Gerda [20]. Goldfish [4]. Gonadal [13]. Grass [6]. Grayling [33]. Growth [9, 33, 24].

Hard [7]. Hard-water [7]. Herefordshire [33]. Herrings [31].

II [29]. indus [19]. infected [28]. Inoculator [34]. Introduced [3]. Ireland [26]. Isles [8, 3]. Israel [22].

Jet [34].

known [27].

L. [30, 14, 12, 29, 9, 13, 33, 18, 21, 2, 19, 4, 23]. Lacépède [3]. Lamprey [13]. Landlocked [13]. Largemouth [3]. Larger [23]. Leuciscus [9, 18, 21, 2]. Literature [11]. Little [27]. Little-known [27]. Llynau [24]. Lough [26]. Lugg [33].

Manx [31]. Mapping [8]. marinus [13]. Marking [34]. Method [28]. Micropterus [3]. Migration [11]. Mixtures [1]. Mizelle [3]. Monogenetic [3]. Mountain [24].

n [20]. near [22]. Neoplasia [17]. North [3]. Northamptonshire [9]. Northern [26]. Note [5, 23].

Observations [26]. Oligochaetes [2]. Operculas [5]. Otoliths [5].

Parasite [12, 29]. Parasites [12, 29]. Petromyzon [13]. Phenol [1]. Photography [5]. Pike [19]. Plaice [20]. Pond [12, 29]. Ponds [22]. Population [18]. Populations [26]. Possible [28]. Preliminary [8, 7]. principalis [3]. Probably [3]. Protein [28]. proteocephalus [18]. Protozoa [20].

Rainboi [32]. Rainbow [1]. Rare [27]. Rearing [24]. Relationship [30, 14]. Relationships [7]. Review [16, 11]. Reviews [10, 15, 25, 35].

Richardson [32]. River [33, 18]. Roach [9]. Rutilus [9].

Salmo [30, 14, 32]. salmoides [3]. Salmon [28, 24]. Salmonid [26]. Salmonids [11]. Scales [5, 24]. Scottish [20]. Scyphidia [20]. Sea [13, 22]. Seas [27]. Seasonal [12]. Serum [28]. Sex [29]. Shulman [20]. Size [30, 29]. Soft [7]. Soft-water [7]. Some [32]. Southern [7]. sp [20]. Spawning [31, 21]. spined [12, 29]. Squalius [9]. Stickleback [12, 29]. Stream [7]. Study [7]. Supply [14]. Survival [30].

Tarns [24]. Teyrn [24]. Three [12, 29]. Three-spined [12, 29]. Thymallus [33]. torulosus [18]. Toxicity [1]. Trematode [3]. Trials [34]. Tributaries [26]. Trichodina [20]. Trout [30, 14, 1, 32]. trutta [30, 14]. Tubificid [2].

U.D.N. [28]. U.D.N.-infected [28]. Upstream [11]. Urocleidus [3]. using [5, 6, 34].

Variation [12, 29]. Variations [19].

Water [22, 7]. Waters [20]. Weed [6]. Willow [9]. Windermere [19].

Yorkshire [12, 29]. Young [20].

Zinc [1].

References

Brown:1969:ATR

[1] V. M. Brown, D. H. M. Jordan, and B. A. Tiller. The acute toxicity to rainbow trout of fluctuating concentrations and mixtures of ammonia, phenol and zinc. Journal of Fish Biology, 1(1):1–9, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1969:TOF

[2] C. R. Kennedy. Tubificid oligochaetes as food of dace Leuciscus leuciscus (L.). Journal of Fish Biology, 1(1):11–15, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maitland:1969:UPM

[3] Peter S. Maitland and Charles E. Price. Urocleidus principalis (Mizelle, 1936), a North American monogenetic trematode new to the British Isles, probably introduced with the largemouth bass Micropterus salmoides (Lacépède, 1802). Journal of Fish Biology, 1(1):17–18, January 1969. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mawdesley-Thomas:1969:FGC

[4] L. E. Mawdesley-Thomas. Furunculosis in the goldfish Carassius auratus (L.). Journal of Fish Biology, 1(1):19–23, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banks:1969:NPF

[5] J. W. Banks and W. Irvine. A note on the photography of fish scales, operculas and otoliths using an enlarger. Journal of Fish Biology, 1(1):25– 26, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cross:1969:AWC

[6] D. G. Cross. Aquatic weed control using grass carp. Journal of Fish Biology, 1(1):27–30, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:1969:PSF

[7] R. H. K. Mann and D. R. O. Orr. A preliminary study of the feeding relationships of fish in a hard-water and a soft-water stream in Southern England. Journal of Fish Biology, 1(1):31–44, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maitland:1969:PAM

[8] P. S. Maitland. A preliminary account of the mapping of the distribution of freshwater fish in the British Isles. Journal of Fish Biology, 1(1):45– 58, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cragg-Hlne:1969:GDL

[9] D. Cragg-Hlne and J. W. Jones. The growth of dace Leuciscus leuciscus (L.), roach Rutilus rutilus (L.) and chub Squalius cephalus (L.) in Willow Brook, Northamptonshire. Journal of Fish Biology, 1(1):59–82, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1969:BRa

[10] Anonymous. Book reviews. Journal of Fish Biology, 1(1):83–84, January 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Banks:1969:RLU

[11] J. W. Banks. A review of the literature on the upstream migration of adult salmonids. Journal of Fish Biology, 1(2):85–136, April 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chappell:1969:PTSa

[12] L. H. Chappell. The parasites of the three-spined stickleback Gasterosteus aculeatus L. from a Yorkshire Pond: I. Seasonal variation of parasite fauna. Journal of Fish Biology, 1(2):137–152, April 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardisty:1969:CGD

[13] M. W. Hardisty. A comparison of gonadal development in the ammocoetes of the landlocked and anadromous forms of the sea lamprey Petromyzon marinus L. Journal of Fish Biology, 1(2):153–166, April 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagenal:1969:RBF

[14] T. B. Bagenal. The relationship between food supply and fecundity in brown trout Salmo trutta L. Journal of Fish Biology, 1(2):167–182, April 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1969:BRb

[15] Anonymous. Book reviews. Journal of Fish Biology, 1(2):183–185, April 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1969:FR

[16] Anonymous. Film review. Journal of Fish Biology, 1(2):186, April 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mawdesley-Thomas:1969:NFB

[17] Lionel E. Mawdesley-Thomas. Neoplasia in fish — a bibliography. Journal of Fish Biology, 1(3):187–207, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1969:PBC

[18] C. R. Kennedy and P. M. Mine. Population biology of the Cestode proteocephalus torulosus (Batsch) in dace Leuciscus leuciscus (L.) of the River Avon. Journal of Fish Biology, 1(3):209–219, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kipling:1969:VFP

[19] C. Kipling and W. E. Frost. Variations in the fecundity of pike Esox indus L. in Windermere. Journal of Fish Biology, 1(3):221–237, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacKenzie:1969:SGA

[20] K. MacKenzie. Scyphidia (Gerda) adunconucleata n. sp. and Trichodina borealis (Dogiel, 1940) Shulman et Shulman — Albova, 1953 (Protozoa Ciliata) from young plaice in Scottish waters. Journal of Fish Biology, 1 (3):239–247, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:1969:SED

[21] Michael Kennedy. Spawning and early development of the dace Leuciscus leuciscus (L.). Journal of Fish Biology, 1(3):249–259, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loya:1969:EFB

[22] Yossef Loya and Lev Flshelson. Ecology of fish breeding in brackish water ponds near the Dead Sea (Israel). Journal of Fish Biology, 1(3):261–278, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sinha:1969:NFL

[23] V. R. P. Sinha. A note on the feeding of larger eels Anguilla anguilla (L.). Journal of Fish Biology, 1(3):279–283, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slnha:1969:SRM

[24] V. R. P. Slnha and H. Evans. Salmon rearing in mountain tarns: The scales and growth of the fish of Llynau Dyrnogydd, Teyrn and Cilan. Journal of Fish Biology, 1(3):285–294, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1969:BRc

[25] Anonymous. Book reviews. Journal of Fish Biology, 1(3):295–296, July 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vickers:1969:OSP

[26] K. U. Vickers. Observations on the salmonid populations of the Lough Erne Tributaries in Northern Ireland. Journal of Fish Biology, 1(4):297– 309, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:1969:RLK

[27] Alwyne Wheeler and R. W. Blacker. Rare and little-known fishes in British Seas in 1966 and 1967. Journal of Fish Biology, 1(4):311–331, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mulcahy:1969:SPC

[28] Maire F. Mulcahy. Serum protein changes in U.D.N.-infected Atlantic salmon: a possible method of diagnosis. Journal of Fish Biology, 1(4): 333–338, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chappell:1969:PTSb

[29] L. H. Chappell. The parasites of the three-spined stickleback Gasterosteus aculeatus L. from a Yorkshire Pond: II. Variation of the parasite fauna with sex and size of fish. Journal of Fish Biology, 1(4):339–347, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagenal:1969:RBE

[30] T. B. Bagenal. Relationship between egg size and fry survival in brown trout Salmo trutta L. Journal of Fish Biology, 1(4):349–353, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bowers:1969:SBM

[31] A. B. Bowers. Spawning beds of Manx autumn herrings. Journal of Fish Biology, 1(4):355–359, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eddy:1969:SEC

[32] F. B. Eddy and R. I. G. Morgan. Some effects of carbon dioxide on the blood of rainboi trout Salmo gairdneri Richardson. Journal of Fish Biology, 1(4):361–372, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hellawell:1969:ADG

[33] J. M. Hellawell. Age determination and growth of the grayling Thymallus thymallus (L.) of the River Lugg, Herefordshire. Journal of Fish Biology, 1(4):373–382, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:1969:FTF

[34] P. J. B. Hart and T. J. Pitcher. Field trials of fish marking using a jet inoculator. Journal of Fish Biology, 1(4):383–385, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:1969:BRd

[35] Anonymous. Book reviews. Journal of Fish Biology, 1(4):387–392, October 1969. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).